# Detecting Throat Cancer from Speech Signals using Machine Learning: A Scoping Literature Review

### v2

@@ 1.1 Background
Identifying cancers is challenging and highly dependent on timely symptom recognition. Common early symptoms of throat cancer include changes in voice, pain or difficulty swallowing, and lumps in the neck [6].
@@
There are many benign voice disorders that have overlapping symptoms with throat cancer, making them difficult to differentiate without these invasive tests [8, 9]. Once cancer has been ruled out, speech assessment can be used to assist in the diagnosis of voice disorders and to measure the degree of abnormality. Two main protocols are commonly used: GRBAS and the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V). GRBAS is a reproducible clinical assessment of abnormal voices commonly used in clinical practice. Various aspects of the patient’s voice are scored on a scale of zero to three, where zero is normal, one is mild, two is moderate, and three is severe, across five different vocal elements: grade, roughness, breathiness, asthenia, and strain [10]. CAPE-V is an assessment tool which requires patients to perform three speech tasks. Six speech features are assessed for each task - overall severity, roughness, breathiness, strain, pitch, and loudness. Each feature is rated on a scale from 0 (normal) to 100 (severe) [11]. These protocols are subjective and require expert knowledge.
With healthcare systems being under immense pressure, it’s important that resources are correctly and most efficiently allocated. In the UK, only approximately 4.3% of patients referred on the urgent suspected cancer pathway for head and neck cancers have a cancer diagnosis; if tools are able to detect non-cancer patients for more efficient referral, some of the load on healthcare systems could be eased [12].
@@
It has been suggested that by using ML or AI, throat cancer may be able to be detected from patient speech [16, 17, 18]. This technology may be able to detect throat cancer sooner and help prioritize those patients who are most at risk while reducing the need for invasive and expensive diagnostic procedures. In this work, we perform a scoping literature review to document and analyse current ML and AI methods for the detection of throat cancer from speech.
This work is structured as follows: Section 2 details the methods used to conduct this search including search terms, inclusion and exclusion criteria, research questions, study selection, and data extraction. Section 3 gives an overview of our results as well as focusing on each of the research questions defined in Section 2. We conclude our work in Section 4.
@@ 2.1 Search Strategy
We derived a set of search terms from the above research questions; the search terms can be seen in Table 1. These terms were discussed and agreed upon with clinical and technical experts.
The literature search was conducted across three databases: Scopus, Web of Science, and PubMed and was run on 08/07/2024. The publication’s title, abstract, and keywords were searched using the terms shown in Table 1. These search terms were chosen to try to cover all of the topics that should be included in articles relevant to this literature search. These topics are cancer, throat, speech, detection, and machine learning. The terms used for cancer, throat, and speech were discussed with an experienced ENT consultant and the terms used for detection and machine learning were discussed with a technical expert to ensure that all relevant terms were included.
@@
tumor* “voice box” spectrograms ml
neoplasm* glottis voice ai
malignan* glottic vocal
“vocal cord” prosody
@@ 2.3 Study Selection
Using the search terms stated in Table 1 across Scopus, Web of Science, and PubMed, a total of 273 articles were found, and 10 additional articles were also identified as being a part of a relevant challenge. We imported all of the articles into Zotero and then used Rayyan to facilitate the removal of duplicates as well as the manual removal of an additional duplicate (n=106) [25, 26]. We screened all article titles and abstracts and excluded ineligible articles (n=138). We then conducted a full article screening and removed any ineligible or inaccessible articles (n=13), resulting in 27 articles being included. Of the articles excluded at full-text screening, the majority (n=9) were excluded because they did not explicitly state the inclusion of cancer patients in their dataset, often looking to precancerous conditions instead. Three articles had the wrong outcome, not classifying pathological speech, and one article was unavailable. Figure 1 shows a flowchart based on the PRISMA framework depicting the process used to identify and screen the articles found in the search [27]. The screening process was carried out by a single researcher.
@@ 2.4 Data Extraction
Figure 2: The typical pipeline used in classifying pathological speech using machine learning. Both preprocessing and feature extraction are optional steps that are not performed by all articles.
@@
Recall=Correctly classified speech of pathology iTotal number of pathology i $\text{Recall}=\frac{\text{Correctly classified speech of pathology i}}{\text{Total number of pathology i speech}}$ (4)
@@ 3.1 Overview
In this section, we provide a detailed overview of the articles included in this study. We grouped the articles based on whether they performed binary or multi-class classification. The extracted data from each article can be seen in Tables 2 and 3 for articles performing binary and multi-class classification, respectively. In both tables, we give the best performance reported in each article in terms of accuracy (Equation 1), and the corresponding specificity (Equation 2), and sensitivity (Equation 3), as those were the most commonly reported metrics. We also report the UAR (Equation 5) for the multi-classification articles, as this was most commonly reported in these articles. We note that two papers performed both binary and multi-class classification and are therefore included in both tables [28, 29]. Of the remaining articles, 12 performed binary classification, and 13 performed multi-class classification. Additionally, one of the articles reviewed is our previous work; despite this, we aim to maintain objectivity in this review [30].
@@ 99.14 X X
Table 2: A summary of the papers using binary classification. MFCC - Mel-frequency cepstral coefficients, STFT - short-term Fourier transform, SVM - support vector machine, XGBoost - extreme gradient boosting, LightGBM - light gradient-boosting machine, ANN - artificial neural network, CNN - convolutional neural network, DT - decision tree, RF - random forest, DNN - deep neural network, MLP - multi-layered perceptron, LVQ - learning vector quantization, GMM - Gaussian mixture models, HMM - hidden Markov model, OSLEM - Online Sequential Extreme Learning Machine, NB - Naive Bayes, LR - logistic regression
@@
Degila et al. [43] X MFCC SVM
@@
Pishgar et al. [47] X MFCC SVM X 88.60 78.23 59.00
@@ 50.66 X X X
Table 3: A summary of the papers using multi-class classification. MFCC - Mel-frequency cepstral coefficients, DNN - deep neural network, NN - neural network, SVM - support vector machine, MIL - multi-instance learning, GMM - Gaussian mixture model, GBT - gradient boosting tree, k-nn - k nearest neighbour, RF - random forest, LR - linear regression, CNN - convolutional neural network
Our review covers articles published between 1996 and 2024, with no date restrictions applied. The distribution of publication years shows a peak of 10 articles published in 2018 due to a challenge that year and an increasing trend in the last three years (2022-2024) with seven articles published. The majority of these articles were conference papers (n=16) compared to journal articles (n=11). However, we do note that the majority of the conference articles (n=10) are from a single challenge.
@@ 3.1.1 Datasets
Table 4 shows the size of the dataset used in each article splitting the datasets into the number of healthy, cancer, and non-cancer pathologies reported by each article. We found that some articles combine cancerous and precancerous conditions, where the number of cancerous and precancerous conditions is made clear we put the number of precancerous samples in the “Non-Cancer Pathology(s)” column. There is a significant outlier in the number of samples used in the articles as Ezzine et al. [35] reports over 6000 samples. We investigated this further and found that in this article, the open-source dataset Saarbruecken Voice Database is used [52]. This dataset contains over 2000 samples ($\sim$2700) and as such the samples referred to by Ezzine et al. cannot be directly from the Saarbruecken Voice Database. We speculate that the authors may have augmented their dataset by splitting audio samples into smaller segments, though this is not explicitly clarified in the article. This discrepancy means that it is impossible to reproduce their results even with the available data. The next largest dataset is fromKwon et al. [18], with 491 total patients. Figure 3 shows the percentage of patients that were healthy, had cancer and had non-cancer pathologies in each article. Gavidia-Ceballos and Hansen [31] has the highest percentage of cancer patients in their dataset (66.7%) with Godino-Llorente and Gomez-Vilda [33] having the lowest percentage (2.2%).
@@
Fang et al. [36] 60 48 354
Kim et al. [16] 45 50 0
Kwon et al. [18] 33 176 282
Wang et al. [17] 0 43 129
Chen et al. [29] 0 40 398
Paterson et al. [30] 38 36 2
Za’im et al. [37] 252 37 93
Kim et al. [28] 155 30 178
Wang et al. [38] 0 60 1940
Verikas et al. [39] 25 22 193
@@ 3.1.2 Preprocessing
Preprocessing is a critical step in data preparation for machine learning, aimed at enhancing the quality of data and improving model performance. In this study, we examined the preprocessing methods reported in the reviewed articles. We found that the majority of articles did not report any preprocessing steps (n=17). Of the articles performing binary classification, only four used preprocessing, each employing different methods. A comparison of the performance of articles using preprocessing versus those not using it revealed no significant difference. For multi-class classification, six articles used preprocessing, while nine did not. Again, no significant difference in results was observed between those that used preprocessing techniques and those that did not.
@@ 3.2 RQ1 - Classification Methods
Many of the articles used multiple classification methods, with an average of 2.2 methods being used per article. The most common are neural networks (of varying types) and support vector machines (SVM).
@@
Figure 4 shows the breakdown of the classification methods used in the articles. For clarity, methods used only once were grouped into the ’other’ category, and all types of neural networks were combined into a single category. Several papers also combined different algorithms into ensemble methods although this is not shown in Figure 4.
Figure 4: The classification methods used across all articles found in this review. The different neural network methods (ANN, BayesNet, CNN, DBN, DNN, LVQ, MLP, NN, RNN, OSELM) were grouped, and any method used only once was put into the other category for ease of visualisation. NN - neural network, SVM - support vector machine, RF - random forest, GMM - Gaussian mixture model, k-nn - k-nearest neighbour.
Figure 5 shows the performance of each of the classification models included in the found articles. This includes all results from articles that use multiple methods.
Figure 5(a) shows the accuracy achieved by each classification method used in articles for binary classification. In the context of binary classification, neural networks and SVMs demonstrated similar performance. LightGBM and XGBoost appeared to perform worse, but due to the limited number of studies using these methods, no definitive conclusion can be drawn.
For multi-class classification, all methods performed similarly in terms of UAR, with neural networks achieving the best results at the extreme. These results can be seen in Figure 5(b).
@@
(b) The UAR achieved in the multi-class classifiers using different classification methods
@@
While these accuracy metrics provide valuable insights, it is important to note that the dataset used plays a crucial role in the performance of each technique. Therefore, comparing articles that use different datasets is challenging.
However, the IEEE FEMH Voice Data Challenge 2018 provides a unique opportunity for direct comparison, as participants were given the same dataset. Participants were provided with 200 samples to be used in training (40 neoplasm, 60 phonotrauma, 50 vocal palsy, and 50 healthy) and models were then tested on an undisclosed test set of 400 samples [41]. Figure 6 shows the UAR found using different methods in the challenge articles. The ten articles submitted to this challenge [44, 46, 45, 40, 48, 42, 41, 49, 47, 43] indicate that all methods perform similarly with overlapping UAR distributions. SVM classifiers exhibited a wider range of performance compared to ensemble methods and neural networks. These results suggest that no single classification method consistently outperforms the others.

### v3

@@ 1.1 Background
Identifying cancers is challenging and highly dependent on timely symptom recognition. Common early symptoms of throat cancer include changes in voice, pain or difficulty swallowing, and lumps in the neck [7]. There are many benign voice disorders that have overlapping symptoms with throat cancer, making them difficult to differentiate without these invasive tests [8, 9]. Once cancer has been ruled out, speech assessment can be used to assist in the diagnosis of voice disorders and to measure the degree of abnormality. Two main protocols are commonly used: GRBAS and the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V). GRBAS is a reproducible clinical assessment of abnormal voices commonly used in clinical practice. Five aspects of the patient’s voice (grade, roughness, breathiness, asthenia, and strain) are scored on a scale of zero to three, where zero is normal, one is mild, two is moderate, and three is severe [10]. CAPE-V is an assessment tool which requires patients to perform three speech tasks, prolonged vowel sounds, reading sentences aloud, and spontaneous speech. Six speech features are assessed for each task - overall severity, roughness, breathiness, strain, pitch, and loudness. Each feature is rated on a scale from 0 (normal) to 100 (severe) [11]. These protocols are subjective and require expert knowledge.
With healthcare systems being under immense pressure, it’s important that resources are correctly and most efficiently allocated. In the UK, only 2.7-4.3% of patients referred on the urgent suspected cancer pathway for head and neck cancers had a cancer diagnosis; if tools are able to detect non-cancer patients for more efficient referral, some of the load on healthcare systems could be eased [12].
@@
It has been suggested that by using ML or AI, throat cancer may be able to be detected from patient speech [16, 17, 18]. This technology may be able to detect throat cancer sooner and help prioritise those patients who are most at risk while reducing the need for invasive and expensive diagnostic procedures. In this work, we perform a scoping literature review to document and analyse current ML and AI methods for the detection of throat cancer from speech.
@@ 2.1 Search Strategy
This literature search was conducted across three databases: Scopus, Web of Science, and PubMed and was run on 14/10/2024. The publication’s title, abstract, and keywords were searched using the terms shown in Table 1. These search terms were chosen to try to cover all topics that should be included in articles relevant to this literature search. These topics are cancer, throat, speech, detection, and machine learning. The terms used for cancer, throat, and speech were discussed with an experienced ENT consultant and the terms used for detection and machine learning were discussed with a technical expert to ensure that all relevant terms were included.
@@
tumor* “voice box” spectrograms “neural network”
neoplasm* glottis voice “deep learning”
malignan* glottic vocal ml
“vocal cord” prosody ai
@@ 2.3 Study Selection
Using the search terms stated in Table 1 across Scopus, Web of Science, and PubMed, a total of 323 articles were found, and ten additional articles were also identified as part of a relevant challenge. All articles were imported into Zotero for reference management [25]. Rayyan was used to facilitate the removal of duplicates (n=127) as well as being used for screening [26]. All article titles and abstracts were screened, and ineligible articles were excluded (n=159). We then conducted a full article screening and removed any ineligible or inaccessible articles (n=10), resulting in 27 articles being included. Of the articles excluded at full-text screening, the majority (n=8) were excluded because they did not explicitly state the inclusion of cancer patients in their dataset, often looking to precancerous conditions instead. One article was the wrong study type, and one article was unavailable. Figure 1 shows a flowchart based on the PRISMA framework depicting the process used to identify and screen the articles found in the search [27]. Two researchers carried out the screening process, and any conflicts were resolved through discussion.
@@ 2.4 Data Extraction
Figure 2: The typical pipeline for classifying pathological speech using machine learning. Both preprocessing and feature extraction are optional steps not performed by all articles.
@@
$R_{i}=\frac{\text{Correctly classified speech of pathology i}}{\text{Total number of pathology i speech samples}}$ (4)
@@ 3.1 Overview
In this section, we provide a detailed overview of the articles included in this study. We grouped the articles based on whether they performed binary or multi-class classification. The extracted data from each article can be seen in Tables 2 and LABEL:tab:MultiResults for articles performing binary and multi-class classification, respectively. In both tables, we give the best performance reported in each article in terms of accuracy (Equation 1) and the corresponding specificity (Equation 2), and sensitivity (Equation 3), as those were the most commonly reported metrics. In Table LABEL:tab:MultiResults We also report the UAR (Equation 5) for the multi-classification articles, as this was most commonly reported in these articles.
Of the 27 articles included in this review 11 articles perform binary classification, 14 articles perform multi-class classification, and two articles perform both binary and multi-class classification. Additionally, we note that one of the articles reviewed is our previous work; despite this, we aim to maintain objectivity in this review [28].
@@ 99.14 X X
Table 2: A summary of the papers using binary classification. The articles are presented in chronological order. MFCC - Mel-frequency cepstral coefficients, STFT - short-term Fourier transform, SVM - support vector machine, XGBoost - extreme gradient boosting, LightGBM - light gradient-boosting machine, ANN - artificial neural network, CNN - convolutional neural network, DT - decision tree, RF - random forest, DNN - deep neural network, MLP - multi-layered perceptron, LVQ - learning vector quantization, GMM - Gaussian mixture models, HMM - hidden Markov model, OSLEM - Online Sequential Extreme Learning Machine, NB - Naive Bayes, LR - logistic regression
@@
Degila et al. [42] X MFCC
Spectral features
SVM
@@
Pishgar et al. [46] X MFCC SVM
XGBoost
LSTM-FCN
X 88.60 78.23 59.00
@@ 51.38 X X X
Song et al. [51] X MFCC
Acoustic features
Demographic data
ResNet 99.69 100 100 X
@@
Table 3: A summary of the papers using multi-class classification. The articles are presented in chronological order. MFCC - Mel-frequency cepstral coefficients, DNN - deep neural network, NN - neural network, SVM - support vector machine, MIL - multi-instance learning, GMM - Gaussian mixture model, GBT - gradient boosting tree, k-nn - k nearest neighbour, RF - random forest, LR - linear regression, CNN - convolutional neural network, LSTM-FCN - long short term memory fully convolutional network
Our review covers articles published between 1996 and 2024, with no date restrictions applied. The distribution of publication years shows a peak of 10 articles published in 2018 due to a challenge that year and an increasing trend in the last three years (2022-2024) with eight articles published. The majority of these articles were conference papers (n=16) compared to journal articles (n=11). However, we do note that the majority of the conference articles (n=10) are from a single challenge.
Tables 2 and LABEL:tab:MultiResults show the range of results achieved in the articles found during this search. In Table 2 accuracy ranges from 80-99%, in Table LABEL:tab:MultiResults accuracy ranges from 50-99% with UAR ranging from 59-81%. There are two articles of note in these tables which achieve exceptional results. In Table 2 it can be seen that Fang et al. [33] achieves an accuracy score of 99.14% on a binary classification task. This, however, was achieved on an external validation which, upon further inspection, does not contain any cancer patients [52]. The best result on the cross-validation was 94.26% on male patients and 90.52% on female patients.
In Table LABEL:tab:MultiResults, we see that Song et al. [51] has exceptional results when compared to the other articles. However, there is a discrepancy in the number of patients in the dataset for each class and the number reported in the confusion matrix (257 patients in the dataset, 977 samples in the confusion matrix). It is unclear from the article why these numbers differ; it is possible that patients provided multiple recordings or audio splitting was conducted. The results presented are calculated from five-fold cross-validation, and as such, we hypothesise that data leakage may have occurred if care had not been taken to ensure that multiple recordings from the same patient had not been split across the validation folds.
During data extraction, all results presented in each article were extracted. The majority of articles present more than one result, with only three articles presenting one result [41, 40, 39]. Chen et al. [34] presents the most results with 44 models being evaluated. The average number of results presented is nine.
We also note that the articles are not consistent in their evaluation methods. Most articles used a single evaluation method, with ten only evaluating on a holdout test set and eight evaluating only on cross-validation. Some articles however, use a mix of evaluation methods, four articles evaluate using cross-validation and a holdout test set, two use cross-validation and an external test set, and one uses cross-validtaion, holdout, and external test sets. It was not clear how the models were evaluated in two of the articles [29, 31]. Figure 3 shows the performance difference between the models evaluated on cross-validation, holdout, and external test sets. All results presented in the articles are considered when creating this figure. It can be seen that when evaluation is performed on cross-validation, the average performance is higher than in articles where evaluation is performed on a holdout test set. The results on external test sets are higher than expected. However, this is largely affected by the results presented by Fang et al. [33] as discussed above, this may be skewed due to discrepancies in the pathologies present in the external dataset.
Figure 3: A comparison of the results obtained when evaluating using cross-validation, holdout, and external test sets. The bar represents the average accuracy, and the error bars are the minimum and maximum values.
@@ 3.1.1 Datasets
Table 4 shows the size of the dataset used in each article splitting the datasets into the number of healthy, cancer, and non-cancer pathologies reported by each article. We found that some articles combine cancerous and precancerous conditions, where the number of cancerous and precancerous conditions is made clear, the number of precancerous samples is added to the “Non-Cancer Pathology(s)” column.
Ben Aicha and Ezzine [31] are unclear in the number of patients in their dataset, simply stating that there are over 100 healthy participants without giving any specific count. There is a significant outlier in the number of samples used in the articles as Ezzine et al. [32] reports over 6000 samples. We investigated this further and found that in this article, the open-source dataset Saarbruecken Voice Database is used [53]. This dataset contains over 2000 samples ($\sim$2700), and as such, the samples referred to by Ezzine et al. cannot be directly from the Saarbruecken Voice Database. We speculate that the authors may have augmented their dataset by splitting audio samples into smaller segments, though this is not explicitly clarified in the article. This discrepancy means that it is impossible to reproduce their results even with the available data.
The next largest dataset is from Wang et al. [37], with 2000 total patients. Figure 4 shows the percentage of patients that were healthy, had cancer and had non-cancer pathologies in each article. Gavidia-Ceballos and Hansen [29] has the highest percentage of cancer patients in their dataset (66.7%) with Godino-Llorente and Gomez-Vilda [30] having the lowest percentage (2.2%).
@@
Kwon et al. [18] 33 176 282
Wang et al. [17] 0 43 129
Chen et al. [34] 0 40 398
Paterson et al. [28] 38 36 2
Song et al. [51] 150 22 85
Za’im et al. [35] 252 37 93
Kim et al. [36] 155 30 178
Wang et al. [37] 0 60 1940
@@ 3.1.2 Preprocessing
Preprocessing is a critical step in data preparation for machine learning, aimed at enhancing the quality of data and improving model performance. In this study, we examined the preprocessing methods reported in the reviewed articles. We found that the majority of articles did not report any preprocessing steps (n=16). Of the articles performing binary classification, only four used preprocessing, each employing different methods. For multi-class classification, six articles used preprocessing, while nine did not.
@@
Figure 5 shows the difference in accuracy when signals were and weren’t preprocessed. It can be seen that the mean values are similar. However, the minimum and maximum values are more extreme when signals are not preprocessed. This may indicate that preprocessing does improve model robustness.
Figure 5: A comparison of the results obtained when signals were preprocessed and when they were not. The bar represents the average accuracy, and the error bars are the minimum and maximum values.
@@ 3.2 RQ1 - Classification Methods
An average of 1.9 model types were used in each article with Pham et al. [47] using the most model types (SVM, random forest, k-nn, gradient boosting, and ensemble). Figure 6 shows the breakdown of the classification methods used in the articles. For clarity, methods have been categorised, although the number of specific methods can still be seen in the stacked bar chat. The most common type of models used are neural networks, followed by support vector machines (SVM).
Figure 6: The classification methods used across all articles found in this review. Models have been categoriesed for ease of understanding.
@@
Figure 7(a) shows the accuracy achieved by each classification method used in articles for binary classification. While neural networks are the most commonly used algorithm, their performance varies greatly, having both the lowest and highest accuracy. Gaussian mixture models have the highest average performance. This figure shows that there is no one model that outperforms the rest of the models.
Figure 7(b) shows the accuracy achieved by each classification method used in articles for multi-class classification. For multi-class classification, all methods performed similarly in terms of accuracy, with neural networks again varying the most, having both the highest and lowest accuracy. k-nn, random forest and gradient boosting all have very similar performance. However, each of these methods is only used once, and so it is difficult to say with any significance that they performed better than the other algorithms. LightGBM has the best average accuracy with ensamble methods close behind. Ensemble methods involve combining multiple classification algorithms with the hopes of enhancing the strengths of each method, this seems to be an effective method from the results seen in Figure 7(b).
@@
(b) The accuracy achieved in the multi-class classifiers using different classification methods
@@
While these accuracy metrics provide valuable insights, it is important to note that the dataset used plays a crucial role in the performance of each technique. Therefore, comparing articles that use different datasets is challenging. However, the IEEE FEMH Voice Data Challenge 2018 provides a unique opportunity for direct comparison, as participants were given the same dataset. Participants were provided with 200 samples to be used in training (40 neoplasm, 60 phonotrauma, 50 vocal palsy, and 50 healthy) and models were then tested on an undisclosed test set of 400 samples [40]. Figure 8 shows the UAR found using different methods in the challenge articles. UAR is used in this case as it is the most commonly reported metric amongst the challenge articles. Ten articles were submitted to this challenge [43, 45, 44, 39, 47, 41, 40, 48, 46, 42].
Only six of the challenge articles report results on a holdout test set [39, 42, 43, 44, 45, 48]. Of the methods reported on a holdout test set (Figure 8(b)), the random forest performs best, although there is only one instance. Neural networks have the second-best average performance but a very wide range, having the single best and worst performance. When evaluated on cross-validation, gradient boosting has a much higher accuracy than the other methods presented. However, this is a single result. Neural networks have the worst average performance, although their performance is the most wide-ranging of all the algorithms. This is a surprising result considering the performance of neural networks when evaluated on a holdout test set. Ensemble methods do well when evaluated using cross-validation but very poorly when evaluated on a holdout test set, this is possibly due to overfitting.
(a) The UAR achieved in the IEEE FEMH Voice Data Challenge 2018 evaluated on cross-validation using different classification methods
(b) The UAR achieved in the IEEE FEMH Voice Data Challenge 2018 evaluated on a holdout test set using different classification methods